Good evening, I would like to ask you from this problem. Please Help me in this.
1)The circumference of front wheel of a cart is 30 ft long and that of the back wheel is 36 ft long. What is the distance travelled by the cart, when the front wheel has done five more revolutions than the rear wheel.

From Yours poor student
E'ma modi

Answer by checkley71(8403)

36*X=30(X+5) OR 36X=30X+150 OR 6X=150 OR X=150/6 OR X=25 REVOLUTIOND FOR THE REAR WHEEL AND 30 REVOLUTIONS FOR THE FRONT WHEEL
36*25=30*30 OR 900=900